Find an equation in slope-intercept form of the line satisfying the specified conditions. Through (-6,-9) perpendicular to -7x-5y=-3 This Problem Really Confuses Me Help Please
Answer by Cromlix(4381)

Sort your equation out to y = mx + c
-7x-5y=-3
-5y = + 7x - 3
5y = - 7x + 3
y = -7/5x + 3/5
Lines that are perpendicular to
one another have slopes(m) that
multiply together to give -1
So, m1 * m2 = - 1
The line above has a slope of -7/5
So, the line perpendicular to it
has a slope of 5/7
-7/5 * 5/7 = -1
Using the equation:
y - b = m(x - a)
Where a = -6, b = -9 and m (slope)= 5/7
y -(-9)= 5/7(x -(-6))
y + 9 = 5/7(x + 6)
y + 9 = 5/7x + 30/7
y = 5/7x + 30/7 - 63/7 (- 9)
y = 5/7x - 33/7
OR multiply thro by 7
7y = 5x - 33
Hope this helps.
:-)